# JWST Observations of the Double Nucleus in NGC 4486B: Possible Evidence for a Recent Binary SMBH Merger and Recoil

### v1

@@ 6 Summary and Conclusion
Although the double nucleus in NGC 4486B was first discovered nearly 30 years ago (Lauer.1996) there have been few studies of its origin and dynamical structure. In analogy with the double nucleus in M31, Lauer.1996 proposed that it was an apsidally aligned eccentric nuclear disk surrounding an SMBH. Two striking differences between M31 and NGC 4486B however are: (a) in M31 the SMBH (located at P3) is at the galactic center while the SMBH in NGC 4486B is 6pc away from the center (defined by the large scale isophotes) and possibly has a velocity offset of $\sim 16\,\mathrm{km\ s}^{-1}$ relative to stars on the opposite side of the galaxy; (b) M31 has a central stellar cusp and a relatively young star cluster coincident with the SMBH while NGC 4486B has a flat core $\gamma=0.14$ with a core radius of $\sim 20$ pc. In Paper 1 (Behzad.2025) we used JWST-NIRSpec IFU spectroscopy and archival HST photometry combined with orbit-superposition Schwarzschild dynamical modeling to confirm the presence of an SMBH of ${\cal M}_{BH}=3.6\pm 10^{8}\mbox{$M_{\odot}$}$. Despite the fact that the Schwarzschild model is axisymmetric and consequently does not reproduce the asymmetric photometry and kinematics, the distribution function provided by the model allows us to examine the origin of the double nucleus and the offset SMBH in NGC 4486B for the first time.
@@
In this work, we focus on the possibility that the double nucleus in NGC 4486B is due to an END similar to that in M31 (Tremaine.1995). We compare the orbital properties of the Schwarzschild model with predicted properties of ENDs that are produced when an SMBH surrounded by a circular nuclear disk experiences a GW recoil kick. These simulations show that a GW recoil following SMBH coalescence can transform an initially circular stellar disk into an eccentric, apsidally aligned structure with a significant retrograde population. Previous studies of surface brightness and kinematic distributions of simulated ENDs match many properties of the observed double nucleus in NGC 4486B: the fainter peak and the velocity dispersion peak coincide and are the location of the SMBH, the asymmetric line-of-sight velocity field, are all consistent with predictions for nearly edge-on ENDs. This interpretation is further reinforced by the finding that our best-fitting Schwarzschild models require a significant fraction of retrograde orbits in the outer nuclear region. Retrograde orbits are not only a natural dynamical consequence of a recoiling SMBH but also a key ingredient in the long-term stability of ENDs. The presence of the central core in this galaxy is additional evidence that it has experienced a binary-black hole mediated “cusp scouring” prior to the binary black hole merger that likely gave the resulting SMBH the kick which produced the END.
@@
From the observed properties of the double nucleus (semi-major axis $\sim 6$ pc) and black hole mass we estimate that the SMBH experienced a recoil kick of $\sim 340\,\mathrm{km\ s}^{-1}$. We adapt the formalism of (Akiba.2025) to estimate the recoil kick magnitude as a function of the pre-merger black hole mass ratio, $q=m_{2}/M_{1}$. Marginalizing over the spin distribution of the two BHs (which is assumed to be isotropic in direction and uniform in scaled magnitude between [0,1]), we estimate that for a recoil kick of $340\,\mathrm{km\ s}^{-1}$, the pre-merger mass ratio was: $q\sim 0.24\pm 0.09$ implying that the mass of the secondary was $m_{2}\sim 7\pm 2.3\times 10^{7}\mbox{$M_{\odot}$}$.
@@
To explore this scenario further, we constructed $N$-body realizations of our best-fit Schwarzschild model from Paper 1 and carried out numerous experiments, which show that a recoiling SMBH, kicked with a velocity of $0.2-0.75v_{\rm esc}\sim 154-576\,\mathrm{km\ s}^{-1}$, oscillates and sinks back to the galactic center within $\sim$10-80 Myr. For our estimated recoil kick of $340\,\mathrm{km\ s}^{-1}$, the timescale to sink to the center is expected to be less than 30 Myr, implying that if the position and velocity offsets of the SMBH are to be attributed to a GW recoil kick, the binary black hole merger in NGC 4486B must have occurred quite recently. The simulations rule out “core stalling” and “dynamical buoyancy” scenarios in which the failure of dynamical friction in a constant density core would allow the SMBH to oscillate for extended periods (Dattathri2025b). This discrepancy is likely due to the overmassive nature of the SMBH in NGC 4486B (${\cal M}_{BH}/M_{*}\sim 4-10$%), while most numerical simulations of this problem assume ${\cal M}_{BH}/M_{*}\sim 0.1-0.2$%, based on expectations of SMBH scaling relations.
@@
Our inference that the recoil event occurred recently presents an apparent paradox. NGC 4486B lies near the dense center of the Virgo Cluster and exhibits a uniformly old, dynamically relaxed stellar population. The absence of tidal features and the uniformly old stars imply that the $\sim 1:4-1:5$ major merger that introduced the secondary black hole must have taken place several gigayears ago, likely when the progenitor galaxies were still in the cluster outskirts.
@@
The implications of this finding extend beyond a single galaxy. If compact ellipticals with double nuclei arising from ENDs like NGC 4486B retain signatures of relatively recent SMBH mergers, they may provide valuable laboratories for studying post-merger dynamics on resolvable scales. Non-equilibrium dynamical models similar to those constructed for M31 (Peiris.2003; Bacon.2001; Sambhus.2002; BrownMagorrian.2013) fitted to the excellent NIRSpec-IFU kinematic maps of NGC 4486B and $N$-body models of the resulting END, embedded inside the rotating elliptical host are needed to more precisely determine the time since the merger event and the expected survival time of the END. NGC 4486B serves not only as a rare nearby case study, but also as a cornerstone for linking high-resolution observations of galactic centers to GW astrophysics.
We are grateful to Tod Lauer for kindly providing the deconvolved F [^5] W image from (Lauer.1996). We thank Eugene Vasiliev for useful discussions. BT and MV acknowledge funding from Space Telescope Science Institute awards: JWST-GO-02567.002-A and HST-GO-16882.002-A. M.A.T. and S.T. acknowledge the support of the Canadian Space Agency (CSA) [22JWGO1-07]. This work was supported by Tamkeen under the NYU Abu Dhabi Research Institute grant CASS, and utilized the high-performance computing resources at New York University Abu Dhabi. CL acknowledges support from the National Natural Science Foundation of China (NSFC, Grant No. 12173025), National Key R&D Program of China (2023YFA1607800, 2023YFA1607804), 111 project (No. B20019), and Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education. B.T. acknowledges support from the Mendel Science Experience (MSE) Postdoctoral Fellowship awarded by Villanova University.
Figure A1: Best-fit Schwarzschild models with (top) and without (bottom) retrograde orbits, each annotated with the corresponding $\chi^{2}$ values.

### v2

@@ VI Summary and Conclusion
Although the double nucleus in NGC 4486B was first discovered nearly 30 years ago (Lauer et al., 1996) there have been few studies of its origin and dynamical structure. In analogy with the double nucleus in M31, Lauer et al. (1996) proposed that it was an apsidally aligned eccentric nuclear disk surrounding an SMBH. Two striking differences between M31 and NGC 4486B however are: (a) in M31 the SMBH (located at P3) is at the galactic center while the SMBH in NGC 4486B is 6 pc away from the center (defined by the large scale isophotes) and possibly has a velocity offset of $\sim 16\,\mathrm{km\ s}^{-1}$ relative to stars on the opposite side of the galaxy; (b) M31 has a central stellar cusp and a relatively young star cluster coincident with the SMBH while NGC 4486B has a flat core $\gamma=0.14$ with a core radius of $\sim 20$ pc. In Paper 1 (Tahmasebzadeh et al., 2025) we used JWST-NIRSpec IFU spectroscopy and archival HST photometry combined with orbit-superposition Schwarzschild dynamical modeling to confirm the presence of an SMBH of ${\cal M}_{BH}=3.6\pm 10^{8}\mbox{$M_{\odot}$}$. Despite the fact that the Schwarzschild model is axisymmetric and consequently does not reproduce the asymmetric photometry and kinematics, the distribution function provided by the model allows us to examine the origin of the double nucleus and the offset SMBH in NGC 4486B for the first time.
@@
In this work, we focus on the possibility that the double nucleus in NGC 4486B is due to an END similar to that in M31 (Tremaine, 1995). We compare the orbital properties of the Schwarzschild model with predicted properties of ENDs that are produced when an SMBH surrounded by a circular nuclear disk experiences a GW recoil kick. These simulations show that a GW recoil following SMBH coalescence can transform an initially circular stellar disk into an eccentric, apsidally aligned structure with a significant retrograde population. Previous studies of surface brightness and kinematic distributions of simulated ENDs match many properties of the observed double nucleus in NGC 4486B: the fainter peak and the velocity dispersion peak coincide and are the location of the SMBH, the asymmetric line-of-sight velocity field, are all consistent with predictions for nearly edge-on ENDs. This interpretation is further reinforced by the finding that our best-fitting Schwarzschild models require a significant fraction of retrograde orbits in the outer nuclear region. Retrograde orbits are not only a natural dynamical consequence of a recoiling SMBH but also a key ingredient in the long-term stability of ENDs. The presence of the central core in this galaxy is additional evidence that it has experienced a binary-black hole mediated “cusp scouring” prior to the binary black hole merger that likely gave the resulting SMBH the kick which produced the END.
@@
From the observed properties of the double nucleus (semi-major axis $\sim 6$ pc) and black hole mass we estimate that the SMBH experienced a recoil kick of $\sim 340\,\mathrm{km\ s}^{-1}$. We adapt the formalism of (Akiba et al., 2025) to estimate the recoil kick magnitude as a function of the pre-merger black hole mass ratio, $q=m_{2}/M_{1}$. Marginalizing over the spin distribution of the two BHs (which is assumed to be isotropic in direction and uniform in scaled magnitude between [0,1]), we estimate that for a recoil kick of $340\,\mathrm{km\ s}^{-1}$, the pre-merger mass ratio was: $q\sim 0.24\pm 0.09$ implying that the mass of the secondary was $m_{2}\sim 7\pm 2.3\times 10^{7}\mbox{$M_{\odot}$}$.
@@
To explore this scenario further, we constructed $N$-body realizations of our best-fit Schwarzschild model from Paper 1 and carried out numerous experiments, which show that a recoiling SMBH, kicked with a velocity of $0.2-0.75v_{\rm esc}\sim 154-576\,\mathrm{km\ s}^{-1}$, oscillates and sinks back to the galactic center within $\sim$10-80 Myr. For our estimated recoil kick of $340\,\mathrm{km\ s}^{-1}$, the timescale to sink to the center is expected to be less than 30 Myr, implying that if the position and velocity offsets of the SMBH are to be attributed to a GW recoil kick, the binary black hole merger in NGC 4486B must have occurred quite recently. The simulations rule out “core stalling” and “dynamical buoyancy” scenarios in which the failure of dynamical friction in a constant density core would allow the SMBH to oscillate for extended periods (Dattathri et al., 2025a). This discrepancy is likely due to the overmassive nature of the SMBH in NGC 4486B (${\cal M}_{BH}/M_{*}\sim 4-10$%), while most numerical simulations of this problem assume ${\cal M}_{BH}/M_{*}\sim 0.1-0.2$%, based on expectations of SMBH scaling relations.
@@
Our inference that the recoil event occurred recently presents an apparent paradox. NGC 4486B lies near the dense center of the Virgo Cluster and exhibits a uniformly old, dynamically relaxed stellar population. The absence of tidal features and the uniformly old stars imply that the $\sim 1:4-1:5$ major galaxy–galaxy merger that introduced the secondary black hole must have taken place several gigayears ago, likely when the progenitor galaxies were still in the cluster outskirts.
@@
The implications of this finding extend beyond a single galaxy. If compact ellipticals with double nuclei arising from ENDs like NGC 4486B retain signatures of relatively recent SMBH mergers, they may provide valuable laboratories for studying post-merger dynamics on resolvable scales. Non-equilibrium dynamical models similar to those constructed for M31 (Peiris & Tremaine, 2003; Bacon et al., 2001; Sambhus & Sridhar, 2002; Brown & Magorrian, 2013) fitted to the excellent NIRSpec-IFU kinematic maps of NGC 4486B and $N$-body models of the resulting END, embedded inside the rotating elliptical host are needed to more precisely determine the time since the SMBH merger event and the expected survival time of the END. NGC 4486B serves not only as a rare nearby case study, but also as a cornerstone for linking high-resolution observations of galactic centers to GW astrophysics.
We are grateful to Tod Lauer for kindly providing the deconvolved F [^5] W image from (Lauer et al., 1996). We thank Eugene Vasiliev for useful discussions. BT and MV acknowledge funding from Space Telescope Science Institute awards: JWST-GO-02567.002-A and HST-GO-16882.002-A. M.A.T. and S.T. acknowledge the support of the Canadian Space Agency (CSA) [22JWGO1-07]. This work was supported by Tamkeen under the NYU Abu Dhabi Research Institute grant CASS, and utilized the high-performance computing resources at New York University Abu Dhabi. CL acknowledges support from the National Natural Science Foundation of China (NSFC, Grant No. 12173025), National Key R&D Program of China (2023YFA1607800, 2023YFA1607804), 111 project (No. B20019), and Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education. B.T. acknowledges support from the Mendel Science Experience (MSE) Postdoctoral Fellowship awarded by Villanova University. MV and BT acknowledge support from NASA through grant number HST-AR-18142 awarded by the Space Telescope Science Institute, which is operated by AURA, Inc., under NASA contract NAS 5-26555. This work is based on observations made with the JWST and obtained from the Mikulski Archive for Space Telescopes (MAST). The data are available at https://doi.org/10.17909/4rr8-zk52 (catalog doi:10.17909/4rr8-zk52).